OpenStudy (anonymous):

b+c=10/a solve for a

OpenStudy (anonymous):

Multiply "A" to the other side, \[A \times B + C = 10\], then minus C to the other side \[A \times B = 10-C\] and then Divide by B to get it to the other side \[A = 10-C \div B\] or even A = 10 -C ------ B

OpenStudy (anonymous):

a=10/(b+c)